OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

OpenStudy (anonymous):

2x-y=-1 -4x+y=-1 -2x=-2 x=-1 Plug x back into either equation and solve for y.